# Tac-Man: Tactile-Informed Prior-Free Manipulation of Articulated Objects

### v3

@@ Acknowledgment
TABLE A1: Hyper-parameters defined in Secs. -A and -B.
@@
Value $10$ $0.25~{}\mathrm{mm}$ $2~{}\mathrm{mm}$ $3$-$6$ $40~{}\mathrm{cm}$ $60~{}\mathrm{cm}$ $4~{}\mathrm{cm}$ $1.5~{}\mathrm{cm}$ $0.02$
As detailed in Sec. III-B, Tac-Man exploits the compliance of the elastic deformable layer on the gripper to guide manipulation. This deformation provides critical tactile feedback by informing the gripper-object contact, which Tac-Man utilizes to adjust the gripper to maintain stable contact. Following previous studies [85], we employ rigid-body simulations to efficiently and rigorously evaluate our method on a wide range of objects. Despite their limitation in handling deformable materials, we do not use soft-body materials due to much lower speed and higher computational demands. To closely replicate the tactile feedback mechanism under these constraints in the rigid-body simulation, we devised specific adaptations to estimate the behavior of a deformable layer, as detailed below. ()$\prescript{\mathrm{sim}}{}{(}\cdot)$ denotes simulated variables.
@@
The discrepancy between $\prescript{\mathrm{sim}}{}{\mathcal{C}_{1}}$ and $\prescript{\mathrm{sim}}{}{\mathcal{C}_{i}}$ in the gripper’s frame quantifies the contact deviation under the non-sliding stable contact conditions (Eq. 2). Based on this information, the iterative process between the execution and recovery stages in the simulation adheres to the methodology detailed in Sec. III and utilizes the same hyper-parameters as in the real-world experiments (Tab. I).
@@
In rigid body simulations, we cannot simulate the elastic deformable layer, which causes misaligned dynamics compared to real-world scenarios. The fingers’ movement directly applies frictional forces to the object handle to move it, rather than through the contact between the layer and the handle. This can lead to unrealistic movements of the object part during the recovery stage, where Eq. 6 does not guarantee a successful recovery. To address this discrepancy, we have empirically increased the damping within the joint settings during this stage to emulate the effect of the deformable layer. This modification suppresses the unintended object movement, aligning the simulated interactions more closely with those observed in real-world conditions. Although these adaptations do not perfectly mirror real-world trajectories, they provide a reasonable framework to assess how our algorithm tracks and maintains contact stability, thereby validating our approach.
@@
Key hyper-parameters play a significant role in our simulation studies, and their values are determined based on empirical observations and testing, which we report in Tab. A1 to ensure transparency and reproducibility of the results.
@@
Our simulation setup incorporates playboards consisting of a baseboard (white) and a toy train (red), and the train is restricted to a predetermined trajectory (grey). To achieve this, we constructed a groove on a cuboid board and set up a distance joint between the board and the train, limiting its movement to within the plane of the board. Two cylindrical locating pins beneath the train’s collision model are inserted into the groove to guide the train along the trajectory. The trajectories are sampled as Bézier curves, $B(s),s\in[0,1]$, with $n_{\mathrm{ctrl}}$ control points uniformly sampled and scaled to fit an $H\mathrm{cm}\times W\mathrm{cm}$ area, resulting in $(n_{\mathrm{ctrl}}-1)$-order curves. Any self-intersecting curves, taking into account a padding of $w$, are discarded. The groove is carved into a $(H+2w)~{}\mathrm{cm}\times(W+2w)~{}\mathrm{cm}$ board following the curve. The initial position of the train is set at $B(\eta)$ with a small positive $\eta$ to prevent initial pin-border collisions. The hyper-parameters for this setup are listed in Tab. A1.

### v4

@@ Acknowledgment
TABLE A1: Hyper-parameters defined in Secs.˜-A and -B.
@@
Value $10$ $0.25~\mathrm{mm}$ $2~\mathrm{mm}$ 33-66 $40~\mathrm{cm}$ $60~\mathrm{cm}$ $4~\mathrm{cm}$ $1.5~\mathrm{cm}$ $0.02$
[…]
@@
The discrepancy between $\prescript{\mathrm{sim}}{}{\mathcal{C}_{1}}$ and $\prescript{\mathrm{sim}}{}{\mathcal{C}_{i}}$ in the gripper’s frame quantifies the contact deviation under the non-sliding stable contact conditions (Eq.˜2). Based on this information, the iterative process between the execution and recovery stages in the simulation adheres to the methodology detailed in Sec.˜III and utilizes the same hyper-parameters as in the real-world experiments (Tab.˜I).
@@
In rigid body simulations, we cannot simulate the elastic deformable layer, which causes misaligned dynamics compared to real-world scenarios. The fingers’ movement directly applies frictional forces to the object handle to move it, rather than through the contact between the layer and the handle. This can lead to unrealistic movements of the object part during the recovery stage, where Eq.˜6 does not guarantee a successful recovery. To address this discrepancy, we have empirically increased the damping within the joint settings during this stage to emulate the effect of the deformable layer. This modification suppresses the unintended object movement, aligning the simulated interactions more closely with those observed in real-world conditions. Although these adaptations do not perfectly mirror real-world trajectories, they provide a reasonable framework to assess how our algorithm tracks and maintains contact stability, thereby validating our approach.
@@
Key hyper-parameters play a significant role in our simulation studies, and their values are determined based on empirical observations and testing, which we report in Tab.˜A1 to ensure transparency and reproducibility of the results.
@@
Our simulation setup incorporates playboards consisting of a baseboard (white) and a toy train (red), and the train is restricted to a predetermined trajectory (grey). To achieve this, we constructed a groove on a cuboid board and set up a distance joint between the board and the train, limiting its movement to within the plane of the board. Two cylindrical locating pins beneath the train’s collision model are inserted into the groove to guide the train along the trajectory. The trajectories are sampled as Bézier curves, $B(s),s\in[0,1]$, with $n_{\mathrm{ctrl}}$ control points uniformly sampled and scaled to fit an $H\mathrm{cm}\times W\mathrm{cm}$ area, resulting in $(n_{\mathrm{ctrl}}-1)$-order curves. Any self-intersecting curves, taking into account a padding of $w$, are discarded. The groove is carved into a $(H+2w)~\mathrm{cm}\times(W+2w)~\mathrm{cm}$ board following the curve. The initial position of the train is set at $B(\eta)$ with a small positive $\eta$ to prevent initial pin-border collisions. The hyper-parameters for this setup are listed in Tab.˜A1.